News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Friday 22 April 2011 9.00am EST

During the last week Chandra completed the observing schedule as planned.

Chandra passed through the second and third eclipses of the spring season on Apr 17 and 20 respectively, with nominal power and thermal performance. Two real-time procedures were completed on Apr 15; the first to update the on-board ephemeris and the second to uplink an update to the on-board gyro scale-factor and alignment matrix. Real-time procedures were executed on Apr 16 to dump OBC-A memory as a follow-up to the gyro scale-factor and alignment matrix patch. The dump will be used to update the baseline memory image maintained on the ground.

The schedule of targets for the next week is shown below and includes one of the series of the Crab Nebula Target of Opportunity observations, observations of FK Com coordinated with the HST, and an observation of PKS 2155-304 coordinated with XMM-Newton and Suzaku.

Markarian3             ACIS-S/HETG  Apr 25
FKCOMAE                ACIS-S/HETG
Radiation Belts
FKCom                  ACIS-S/HETG  Apr 26
1RXSJ171405.2-20274    ACIS-S
PKS2155-304            ACIS-S/LETG
1RXSJ173916.2-21474    ACIS-S       Apr 27
FKCom                  ACIS-S/HETG
Markarian3             ACIS-S/HETG
CrabNebula             ACIS-S       Apr 28
J082355.0+280621.8     ACIS-S
Radiation Belts
Markarian3             ACIS-S/HETG
FKCOMAE                ACIS-S/HETG  Apr 29
1ES1553+113            HRC-S/LETG
A1914_Field4           ACIS-I       Apr 30
Markarian3             ACIS-S/HETG


All spacecraft subsystems continued to support nominal operations.

Return to Status Reports